OpenStudy (anonymous):

What is the simplified form of 15 times x to the fifth power over 9 times y to the seventh power divided by45 times y-cubed over 5 times x squared ?

OpenStudy (anonymous):

hey can you write the equation plz

OpenStudy (anonymous):

its confusing

OpenStudy (anonymous):

\[\huge\frac{\left(\frac{15x^5}{9y^7}\right)}{\left(\frac{45y^3}{5x^2}\right)}?\]

OpenStudy (anonymous):

Siths has it right, but I don't understand how to solve the problem

OpenStudy (anonymous):



OpenStudy (anonymous):

this is the answer

OpenStudy (anonymous):

Okay, but it's not showing as a possible answer, the possible answers are: 5 x to the seventh over 27 y to the tenth power 15 x to the third power over y to the fourth y to the fourth power over 15 x-cubed or 27 y to the seventh power over 5 x to the seventh power

OpenStudy (anonymous):

simplify it to lowest factor ans is first one

OpenStudy (anonymous):

okay, Thanks!